Hello.
I am currently working me way through a pre-course docket for a royal navy engineering course. I have a GCSE B maths grade (in 1999) and my use of maths has been fairly limited. However, I have been doing quite well in this docket, until now.
I have nipped forward to find the answer to this exerise that follows. I cannot work out how the answer was obtained; I also have no faith in this dockets authority as a number of terrible typos have beset me already.
The equation is thus:
5x+24=3x
That is exactly how it is printed in the docket. The answer, as printed, is apparently -12.
Please, could you explain to me, how this could be? I also know that with an unkown on both sides, it is customary to eliminate x from one side, however, this has still not lead to the answer of -12.
Could this be right? Or could it be the result of any number of typing errors?
Thanks for your time.
Craig Jenkins (panicking)

Found 2 solutions by janakisubramanian, CharlesG2:
Answer by janakisubramanian(2)

Hi,
The given equation is 5x + 24 = 3x
Transpose 3x to LHS
==> 5x + 24 - 3x = 0
==> 5x - 3x + 24 = 0 {rearranging}
==> 2x + 24 = 0 {simplifying like terms}
==> 2x = -24 {transposing 24 to RHS}
==> x = -24/2
==> x = -12
Thus, the given answer is right. i.e. x = -12
Good luck!!
Janaki

Answer by CharlesG2(834)

"Hello.
I am currently working me way through a pre-course docket for a royal navy engineering course. I have a GCSE B maths grade (in 1999) and my use of maths has been fairly limited. However, I have been doing quite well in this docket, until now.
I have nipped forward to find the answer to this exerise that follows. I cannot work out how the answer was obtained; I also have no faith in this dockets authority as a number of terrible typos have beset me already.
The equation is thus:
5x+24=3x
That is exactly how it is printed in the docket. The answer, as printed, is apparently -12.
Please, could you explain to me, how this could be? I also know that with an unkown on both sides, it is customary to eliminate x from one side, however, this has still not lead to the answer of -12.
Could this be right? Or could it be the result of any number of typing errors?
Thanks for your time.
Craig Jenkins (panicking)
"
I sincerely wish you luck and success in your endeavors. I also hope your typo problems get resolved.
5x + 24 = 3x I believe that is what you would like explained.
5x + 24 = 3x
5x - 3x + 24 = 3x - 3x (notice did same operation on both sides)
2x + 24 = 0
2x + 24 - 24 = 0 - 24 (same operation both sides)
2x = -24
x = -12 (divided both sides by 2)
The answer you had of -12 does appear to be correct.
I hope this helps.